Room capacity is 250 m3, air temperature 18 ° C and relative humidity 45%.
How many kilograms of water in the room air and how much air can fit in the air
water as water vapor? The total pressure is 101.3 kPa.
1
Expert's answer
2019-03-12T11:10:05-0400

Solve the first part by determining the absolute humility in "\text{g/m}^3" using the psychrometric tables. For "T=18^\circ\text{C}" find the saturation vapor density ("\rho_{SVD}"): it equals "15.4 \text{ g/m}^3". Hence:


"AH=\frac{RH}{100\%}\cdot \rho_{SVD}=\frac{45\%}{100\%}\cdot 15.4=6.93\text{ g/m}^3."

Since we know the total mass of water in grams per one cubic meter, it is easy to calculate the total mass of water in the room air:


"m=\frac{AH}{1000}\cdot V=\frac{6.93}{1000}\cdot 250=1.73\text{ kg}."



How much vapor the air can contain in these conditions is determined by the saturation vapor density ("\rho_{SVD}") which we have found in the psychrometric tables. It equals "15.4 \text{ g/m}^3", so the air in the room can contain



"m_{\text{max}}=V\cdot\frac{\rho_{SVD}}{1000}=250\cdot\frac{15.4}{1000}=3.85\text{ kg}"

of water.